# Fair Uncertainty Quantification for Depression Prediction

### v2

@@ III-C FUQ for depression prediction
In order to achieve this optimal fairness-aware UQ, it is necessary to transform the optimizing sub-coverage $\beta_{m,s}$ into optimizing the quantile sub-threshold $\hat{r}_{m,s}$ of conformal scores within the group conditional conformal prediction framework. To guarantee overall statistical validity, we initially disregarded sensitive attributes and acquired the quantile $\hat{r}$ of conformal scores for the complete calibration set. For each sensitive-attribute bin, the initial sub-coverage $\beta_{m,s}^{0}$ is computed using $\hat{r}$. $\beta_{m,s}^{0}$ serves as the starting point for the fairness-aware optimization. Although the initial sub-coverage may not be perfectly fair across demographic groups, it still satisfies overall statistical validity, which is the primary advantage of this initialization. If we straightly perform group calibration, with a significantly reduced sample size, the validity guarantee of the CP is also reduced.
@@
After sub-coverage initialization, we need to compare the coverage of each sensitive attribute, i.e., compute the average coverage of $M$ bins. For over-covered groups of sensitive attributes, we need to decrease $\hat{r}_{m,s}$. Conversely, under-covered groups of sensitive attributes need to increase $\hat{r}_{m,s}$ (ideal coverage is $1-\alpha$).
Algorithm 2 Fairness-aware Optimization for Prediction Interval
0:  $\mathcal{D}_{cal}=\{X_{i}^{s},y_{i}\}_{i=1}^{N},s\in[1,S]$: calibration set;
$\beta_{m,s}^{0},m\in[1,M]$: sub-coverage for $m$-th bin with attribute $s$;$\mathcal{Q}(\cdot,\cdot)$: quantile estimator;$B_{1},B_{2},...,B_{M}$: $M$ equal-mass bins; $|B|$: sample sizes of the bin $B$;$r_{m,s}$: conformal scores for whose $y$ belongs to bin $m$ with attribute $s$;$E$: max iteration round.
0:  Conformal threshold $\hat{r}_{m,s}$.
1:  for $e=1$ to $E$ do
2:     Initialize $\beta_{s}^{e}=\frac{1}{M}\sum_{m=1}^{M}\beta_{m,s}^{e-1}$ for all $s\in[1,S]$
3:     $s_{\max}=\arg\max\beta_{s}^{e}$
4:     $s_{\min}=\arg\min\beta_{s}^{e}$
5:     if $\min\beta_{s}^{e}\geq 1-\alpha$ then
6:        break
7:     end if
8:     for $m=1$ to $M$ do
9:        $\nabla\hat{r}_{m}^{-}\leftarrow\frac{\mathcal{Q}(\beta_{m,s_{\max}}^{e},r_{m,s_{\max}})-\mathcal{Q}(\beta_{m,s_{\max}}^{-},r_{m,s_{\max}})}{|B_{m,s_{\max}}|}$
10:        $\nabla\hat{r}_{m}^{+}\leftarrow\frac{\mathcal{Q}(\beta_{m,s_{\min}}^{+},r_{m,s_{\min}})-\mathcal{Q}(\beta_{m,s_{\min}}^{e},r_{m,s_{\min}})}{|B_{m,s_{\min}}|}$
11:     end for
12:     $d=\min\left(\frac{1}{|B_{m,s_{\max}}|},\frac{1}{|B_{m,s_{\min}}|}\right)$
13:     $\nabla\hat{r}^{-}=\max\nabla\hat{r}_{m}^{-}$
14:     $\nabla\hat{r}^{+}=\min\nabla\hat{r}_{m}^{+}$
15:     $\hat{r}_{m,s_{\max}}^{e+1}=\hat{r}_{m,s_{\max}}^{e}-\nabla\hat{r}^{-}\cdot d$
16:     $\hat{r}_{m,s_{\min}}^{e+1}=\hat{r}_{m,s_{\min}}^{e}+\nabla\hat{r}^{+}\cdot d$
17:     Initialize $\beta_{m,s}^{e}=0$ for all $m,s$
18:     for $i=1$ to $N$ do
19:        for $m=1$ to $M$ do
20:           Compute coverage interval for $y_{i}\in B_{m}$ and $s=s^{\prime}$
21:           $C(X_{i}^{s})=\left[\hat{q}_{\alpha/2}(X_{i}^{s})-\hat{r}_{m,s}^{e+1},\hat{q}_{1-\alpha/2}(X_{i}^{s})+\hat{r}_{m,s}^{e+1}\right]$
22:           $\beta_{m,s}^{e}\leftarrow\beta_{m,s}^{e}+\frac{\mathbf{1}\{y_{i}\in B_{m}\cap C(X_{i})\}\cdot\mathbf{1}\{X_{i}^{s}=X_{i}^{s^{\prime}}\}}{|B_{m}|}$
23:        end for
24:     end for
25:  end for
26:  return  $\hat{r}_{m,s}$
@@
As for the updating method, we need to use the slope of $\beta_{m,s}$, instead of the gradient, since we cannot guarantee the differentiable property. Specifically, we can think of the overall validity guarantee as covering $(1-\alpha)*|\mathcal{D}_{cal}|$ samples on the calibration set. Then, the over-coverage group should cover one fewer sample, while the under-coverage group should cover one more sample. And we can represent the effect of covering one less (or one more) sample on the width of the interval by computing the slope of $\beta_{m,s}$. Furthermore, in order to obtain optimal interval widths, it is necessary to ensure that the over-coverage group maximizes the slope when it covers one fewer sample (i.e., decreases the interval width as much as possible), and the under-coverage group minimizes the slope when it covers one more sample (i.e., increases the interval width as little as possible). The natural termination condition for optimization is that both the under-coverage and over-coverage groups achieve the desired coverage, or that the slope of the under-coverage group is greater than that of the over-coverage group, shown in Algorithm 2.
TABLE I: Results of depression prediction on different datastes
@@ IV-A Dataset
MODMA [49]: This is a multimodal depression recognition dataset proposed by Lanzhou University, containing both audio and EEG modalities. The dataset consists of 52 participants (23 major depressive disorder subjects and 29 healthy control subjects), with ages ranging from 18 to 25 years ($<$30: $\geq$30 = 1: 1), and PHQ-9 questionnaire scores as labels. We randomly split the data into training, validation, and test sets using a 3: 1: 1 ratio. The settings are as follows: Hanning window is set to 1024, audio sample rate is 44.1 kHz, hop length is 128, and Mel filters are 64.
@@ IV-B1 Predictive Model
For the predictive model of AVEC 13 & 14, we adopted a C3D network. Following previous studies [50], we adjusted the penultimate fully connected (FC) layer to 64. For the DAIC-WOZ audio predictive model, we designed a 1D CNN architecture inspired by [51]. We utilized the Adam optimizer with a weight decay of 0.0001 and a learning rate of 0.0001. Similarly, for the MODMA audio predictive model, we adopted a network configuration similar to that of the DAIC-WOZ audio model.
For all three networks, the final FC layer consists of 99 units, corresponding to the quantile settings from 0.01 to 0.99 for pinball loss. This design enables the training of a multi-output quantile regression model. We consider the output corresponding to the 50% quantile as the model’s prediction. Alternatively, multiple single-output quantile regression models could also be trained separately. With well training, we can obtain the quantile regression models $\hat{q}_{lo}$, $\hat{q}_{hi}$, and the predictive model $f$.
IV-B2 Trustworthy Methods for Depression Prediction
In regression tasks, CP [52] and CQR [46] can obtain valid prediction intervals. CP does not require retraining of the model and can be utilized for vanilla regression. The disadvantage of CP is that the width of the prediction interval is fixed, which is not suitable for individuals with varying features, such as depressed patients. In contrast, CQR necessitates retraining the original single‑output depression prediction model, converting it into a multi‑output quantile regression model, thereby yielding adaptive prediction distributions for different individuals. This makes the prediction intervals align more closely with the true uncertainty of depression predictions. However, both methods ignore fairness. Although group-wise partitioning can be used to further mitigate disparities across groups, it often results in a reduced sample size within each group, which can compromise the statistical effectiveness of post-hoc methods such as conformal prediction. CFQR [53] considers the DP-fairness of the quantile estimates in CQR. However, applying fairness post-processing using the Wasserstein barycenter may introduce excessive risk [54]. In contrast, BFQR [29] leverages a valid interval optimization strategy to adjust the interval width, achieving improved calibration performance. However, the optimization strategy in BFQR primarily focuses more on overall effectiveness. In this context, our FUQ further refines the optimization strategy by incorporating a fairness-aware optimization framework for prediction interval, thereby achieving a more equitable coverage rate across demographic groups.
TABLE II: Trustworthy predictions of gender group between different methods
Method Dataset PICP(%) MPIW PICP(s = 0) PICP (s = 1) PICP Gap
@@ IV-B2 Trustworthy Methods for Depression Prediction
$\alpha=0.1,s=0$ for female, $s=1$ for male.
TABLE III: Trustworthy predictions of age group between different methods
Method Dataset PICP(%) MPIW PICP(s = 0) PICP (s = 1) PICP Gap
@@ IV-B3 Metrics
Figure 3: Performance of different fairness methods on gender group across various datasets. At the 90 % confidence level ($\alpha=0.1$), our method yields the smallest group-wise coverage gap.
@@
Figure 4: Performance of the various fairness methods on the MOMDA test set across different demographic groups: (a) gender and (b) age. At the 90 % confidence level ($\alpha=0.1$), our method yields the smallest group-wise coverage gap.
@@ IV-C Results
IV-D Limitations and Future Work

### v3

@@ III-B FUQ for Depression Prediction
In order to achieve this optimal fairness-aware UQ, it is necessary to transform the optimizing sub-coverage $\beta_{m,s}$ into optimizing the quantile sub-threshold $\hat{r}_{m,s}$ of conformal scores within the group conditional conformal prediction framework. To guarantee overall statistical validity, we initially disregarded sensitive attributes and acquired the quantile $\hat{r}$ of conformal scores $\beta_{m}$ for the complete calibration set. For each sensitive-attribute bin, the initial sub-coverage $\beta_{m,s}^{0}$ is computed using $\hat{r}_{m,s}=\hat{r}$. $\beta_{m,s}^{0}$ serves as the starting point for the fairness-aware optimization. $\hat{r}_{m,s_{2}}$ (ideal coverage is $1-\alpha$).
$C(X_{i}^{s})=\left[f_{\alpha/2}(X_{i}^{s})-\hat{r}_{m,s},f_{1-\alpha/2}(X_{i}^{s})+\hat{r}_{m,s}\right]$ (6)
$\beta_{m,s}\leftarrow\frac{\mathbf{1}\{y_{i}\in(B_{m}\cap C(X_{i}^{s}))\}}{|B_{m}|}$ (7)
Although the initial sub-coverage may not be perfectly fair across demographic groups, it still satisfies overall statistical validity, which is the primary advantage of this initialization. If we straightly perform group calibration, with a significantly reduced sample size, the validity guarantee of the CP is also reduced.
@@
After sub-coverage initialization, we need to compare the coverage of each sensitive attribute, i.e., compute the average coverage of $M$ bins. For over-covered groups of sensitive attributes, for example $s_{1}$, we need to decrease $\hat{r}_{m,s_{1}}$. Conversely, under-covered groups of sensitive attributes, for example $s_{2}$, need to increase $\hat{r}_{m,s_{2}}$. As for the updating method, we need to use the slope of $\beta_{m,s}$, instead of the gradient, since we cannot guarantee the differentiable property. For m=1,,Mm=1,...,M, we can acquire $\nabla\hat{r}_{m}^{-}$ and $\nabla\hat{r}_{m}^{+}$.
$\nabla\hat{r}_{m}^{-}\leftarrow\frac{\mathcal{Q}(\beta_{m,s_{0}},r_{m,s_{0}})-\mathcal{Q}(\beta_{m,s_{0}}^{-},r_{m,s_{0}})}{|B_{m,s_{0}}|}$ (8)
$\nabla\hat{r}_{m}^{+}\leftarrow\frac{\mathcal{Q}(\beta_{m,s_{1}}^{+},r_{m,s_{1}})-\mathcal{Q}(\beta_{m,s_{1}}^{e},r_{m,s_{1}})}{|B_{m,s_{1}}|}$ (9)
@@
Furthermore, we can think of the overall validity guarantee as covering $(1-\alpha)*|\mathcal{D}_{cal}|$ samples on the calibration set. The over-coverage group should cover one fewer sample, while the under-coverage group should cover one more sample. In order to obtain optimal interval widths, it is necessary to ensure that the over-coverage group maximizes the slope, $\nabla\hat{r}^{-}=\max\nabla\hat{r}_{m}^{-}$, when it covers one fewer sample (i.e., decreases the interval width as much as possible). The under-coverage group minimizes the slope, $\nabla\hat{r}^{+}=\min\nabla\hat{r}_{m}^{+}$, when it covers one more sample (i.e., increases the interval width as little as possible). We can represent the effect of covering one less (or one more) sample on the width of the interval by computing the slope of $\beta_{m,s}$ and acquiring $\hat{r}_{m,s}$.
$\hat{r}_{m,s_{1}}=\hat{r}_{m,s_{1}}-\nabla\hat{r}^{-}$ (10)
$\hat{r}_{m,s_{2}}=\hat{r}_{m,s_{2}}+\nabla\hat{r}^{+}$ (11)
We update $\hat{r}_{m,s}$ using the fixed-point iterations in Eqs. 6-11 before achieving EOC.The natural termination condition for optimization is that both the under-coverage and over-coverage groups achieve the desired coverage, or that the slope of the under-coverage group is greater than that of the over-coverage group.
During evaluation, for the participant $X_{test}^{s}$ with a given attribute $s$, we hypothetically assign the depression score to each severity bin $B_{m}$; for each $B_{m}$, we construct the FUQ interval with the corresponding calibration threshold $\hat{r}_{m,s}$ and enforce feasibility by intersecting with the bin bounds. The overall prediction interval is obtained as their union.
TABLE I: Prediction results on different attribute groups
@@ IV EXPERIMENTS
To validate the effectiveness of FUQ, we conducted experiments on several common datasets, including AVEC [31, 32], DAIC-WOZ [49], and MODMA [50]. For generality, we covered audio and video modalities and relied on methods with public code and pretrained weights (or otherwise easy to reproduce). We then compared FUQ with established trustworthy prediction baselines and further analyzed the fairness in UQ.
@@ IV-A Dataset
MODMA [50]: This is a multimodal depression recognition dataset proposed by Lanzhou University, containing both audio and EEG modalities. The dataset consists of 52 participants (23 major depressive disorder subjects and 29 healthy control subjects), with ages ranging from 18 to 25 years (<<30: $\geq$30 = 1: 1), and PHQ-9 questionnaire scores as labels. In this experiment, we focus on the audio modality and randomly split the data into training, validation, and test sets using a 3: 1: 1 ratio. We extracted participant audio over time frames and utilized log-Mel spectrogram features. The settings are as follows: Hanning window is set to 1024, audio sample rate is 44.1 kHz, hop length is 128, and Mel filters are 64.
@@ IV-B1 Predictive Model
For the predictive model on AVEC 13 & 14, we adopted a C3D network. Following previous studies [52], we adjusted the penultimate fully connected (FC) layer to 64. For the DAIC-WOZ audio dataset, we designed a 1D CNN architecture inspired by [53]. We utilized the Adam optimizer with a weight decay of 0.0001 and a learning rate of 0.0001. Similarly, for the MODMA audio dataset, we adopted a network configuration similar to that of the DAIC-WOZ dataset.
For all three networks, the final FC layer consists of 99 units, corresponding to the quantile settings from 0.01 to 0.99 for pinball loss. This design enables the training of a multi-output quantile regression model. We consider the output corresponding to the 50% quantile as the model’s prediction. Alternatively, multiple single-output quantile regression models could also be trained separately. With well training, we can obtain the quantile regression model $f$.
IV-B2 Trustworthy Depression Prediction
In regression tasks, CP [54] and CQR [47] can obtain valid prediction intervals. CP does not require retraining of the model and can be utilized for vanilla regression. The disadvantage of CP is that the width of the prediction interval is fixed, which is not suitable for individuals with varying features, such as depressed patients. In contrast, CQR necessitates retraining the original single‑output depression prediction model, converting it into a multi‑output quantile regression model, thereby yielding adaptive prediction distributions for different individuals. This makes the prediction intervals align more closely with the true uncertainty of depression predictions. However, both methods ignore fairness. Although group-wise partitioning can be used to further mitigate disparities across groups, it often results in a reduced sample size within each group, which can compromise the statistical effectiveness of post-hoc methods such as conformal prediction. CFQR [55] considers the fairness of the quantile estimates in CQR under the requirement of Demographic Parity. However, applying fairness post-processing using the Wasserstein barycenter may introduce excessive risk [56]. In contrast, BFQR [30] leverages a valid interval optimization strategy to adjust the interval width, achieving improved calibration performance. However, the optimization strategy in BFQR primarily focuses more on overall effectiveness. In this context, our FUQ further refines the optimization strategy by incorporating a fairness-aware optimization framework for prediction intervals, thereby achieving a more equitable coverage rate across demographic groups.
TABLE II: Predictive reliability on gender group for different methods
Method Dataset PICP(%) MPIW PICP(s = 1) PICP (s = 2) PICP Gap $\downarrow$
@@ IV-B2 Trustworthy Depression Prediction
$\alpha=0.1,s=1$ for female, $s=2$ for male.
TABLE III: Predictive reliability on age group for different methods
Method Dataset PICP(%) MPIW PICP(s = 1) PICP (s = 2) PICP Gap $\downarrow$
@@ IV-B3 Metrics
Figure 3: Performance of different methods on gender group across various datasets. At the 90 % confidence level ($\alpha=0.1$), our method yields the smallest group-wise coverage gap.
@@
Figure 4: Performance of various methods on the MOMDA test set across different demographic groups: (a) gender and (b) age. At the 90 % confidence level ($\alpha=0.1$), our method yields the smallest group-wise coverage gap.
@@ IV-C Results
IV-D Parameter Analysis
$M$ plays a pivotal role in FUQ’s fairness performance. Increasing $M$ decreases the sample sizes of each bin, degrading the reliability of UQ. Decreasing $M$ weakens conditional approximation, drifting toward marginal coverage and producing wider, less informative intervals (under equal confidence, narrower intervals are considered better). In our experiments, we split the calibration set into two disjoint parts to select the optimal hyperparameter $M$. We then fix $M$ and calculate the $\hat{r}_{m,s}$ using the entire calibration set. We do not pursue fairness at all costs; rather, we seek group FUQ while maintaining an overall coverage guarantee. As shown in Figs. 5 and 6, the optimal value of the parameter ($M=4$) can be obtained by tuning on the validation set.
IV-E Limitations and Future Work
@@
Figure 5: PICP and MPIW with varying $M$ on AVEC 2014.
Figure 6: PICP Gap with varying $M$ on AVEC 2014.